OpenStudy (anonymous):

What is the solution of 4x^2+1>=4x (answer should be in [ or ) possible an "and" or an "or"

OpenStudy (anonymous):

When dealing with quadratics, it helps to make it equal zero so we first subtract 4x from both sides: 4x^2+1 >= 4x 4x^2-4x+1 >= 0 4(x^2-x)+1 >= 0 (Factor out the 4) 4(x-0.5)^2 >= 0 (Write in completed square format) (x-0.5)^2 >= 0 (Divide both sides by 4) Taking the square root of both sides gives us two answers: +(x-0.5) >= 0 So x >= 0.5 -(x-0.5) >= 0 So x <= 0.5 So we know x >= 0.5 OR x <= 0.5. Therefore x is any real number :)